a research report that the average iq level of students in philiipines science highschool is 110 a sample of 20 students has a mean iq level of 106 woth a standard deviaton of 9


1
Expert's answer
2021-07-16T02:21:23-0400

"\mu = 110 \\

n=20 \\

\bar{x}=106 \\

s=9 \\

H_0: \mu= 110 \\

H_1: \mu≠110"

When n<30 and "\sigma" is unknown a one sample t-test is used.

"t= \frac{\bar{x}- \mu}{s/ \sqrt{n}} \\

t = \frac{106-110}{9 / \sqrt{20}}= -1.987 \\

df=n-1=20-1=19 \\"

Let use α=0.05

Two-tailed test

"t_{\u03b1\/2,df} = t_{0.025, 19}=2.093"

If test statistics "|t| > t_{\u03b1\/2,n-1}" , you have sufficient evidence to reject H0.

"|t|=1.987 < t_{0.025, 11}=2.093"

Accept H0.

There is enough evidence to conclude that the average iq level of students in philippines science highschool is 110 at the 0.05 significance level.